OpenStudy (anonymous):

simplify by rationalizing the denominator 7/√21

jigglypuff314 (jigglypuff314):

to rationalize a denominator understand that raticals with same base multiplied by each other cancel each other out... like √3 * √3 = 3 and √5 * √5 = 5 so to rationalize the denominator of 7/√21 multiply by a value equivalent to 1 like (√21/√21) so \[\frac{ 7 }{ \sqrt{21} } = \frac{ 7 }{ \sqrt{21} }*\frac{ \sqrt{21} }{ \sqrt{21} } = \frac{ 7*\sqrt{21} }{ 21 }\]

jigglypuff314 (jigglypuff314):

and the 7/21 part of it can be simplified down more :)